factoring 9(y-1)^2-25.
a) I'm sorry if I put this question in the wrong spot. I didn't know where to put it.
b) I know the answer is: (3y-8)(3y+2), but I don't understand how to get that.

Answer by ewatrrr(24785)

9(y-1)^2-25 = 9(y^2 - 2y + 1) - 25 = 9y^2 - 18y -16
9y^2 - 18y -16 = (3y -8)(3y+2)
F First terms
O Outside terms 6y
I Inside terms -24y
L Last terms